OpenStudy (anonymous):

solve for x and y. 3x-8i=12-(12y)i please help

OpenStudy (anonymous):

For two complex numbers to be equal to each other, both the real parts and the imaginary parts of the number have to be equal. Therefore, you can separate that into 2 equations: Real(3x-81) = Real(12-12yi) <=> 3x = 12 <=> x = 4 Imaginary(3x-8i) = Imaginary(12-12yi) <=> -8 = -12y <=> y = 12/8 = 3/2

OpenStudy (anonymous):

Little mistake there, should be 8/12 = 2/3, and not 12/8

OpenStudy (anonymous):

unh where did u get Real(3x-81)

OpenStudy (anonymous):

should be 8*i , not 81, sorry

OpenStudy (anonymous):

ok,i got it